To compare the results of boys and girls in a class, a special test was given to 50 boys who averaged 67.4 with Standard deviation of 5, and 50 girls averaged 62.8 with Standard deviation of 4.6.

a) Test, at alpha=0.05, whether the difference is significant or not.

b) Test if the difference is 3.0.


1
Expert's answer
2022-01-26T17:50:38-0500

Boys

"n_1=50\\\bar x=67.4\\s_1=5"

Girls

"n_2=50\\\bar x_2=62.8\\s_2=4.6"

"a)"

We test the following hypotheses,

"H_0:\mu_1-\mu_2=0\\vs\\H_1:\mu_1-\mu_2\not=0"

We apply t distribution to perform this test as follows,.

The test statistic is given as,

"t_c={(\bar x_1-\bar x_2)-(\mu_1-\mu_2)\over \sqrt{{s_1^2\over n_1}+{s_2^2\over n_2}}}={(67.4-62.8)-0\over\sqrt{{25\over50}+{21.16\over 50}}}={4.6\over\sqrt{0.9232}}={4.6\over0.96083297}=4.7875"

"t_c" is compared with the table value at "\alpha=0.05" with "n_1+n_2-2=50+50-2=98" degrees of freedom.

The table value is,

"t_{{0.05\over2},98}=t_{0.025,98}=1.984467"

The null hypothesis is rejected if "|t_c|\gt t_{0.025,98}."

Now,

"|t_c|=4.7875\gt t_{0.025,98}=1.984467," and we reject the null hypothesis and conclude that there is sufficient evidence to show that the difference in means between boys and girls is significant at 5% level of significance.


"b)"

The hypothesis tested are,

"H_0:\mu_1-\mu_2=3\\vs\\H_1:\mu_1-\mu_2\not=3"

"t_c={(\bar x_1-\bar x_2)-(\mu_1-\mu_2)\over \sqrt{{s_1^2\over n_1}+{s_2^2\over n_2}}}=={(67.4-62.8)-3\over\sqrt{{25\over50}+{21.16\over 50}}}={(4.6-3)\over\sqrt{0.9232}}={1.6\over 0.96083297}=1.6652"

"t_c" is compared with the table value at "\alpha=0.05" with "n_1+n_2-2=50+50-2=98" degrees of freedom.

The table value is,

"t_{{0.05\over2},98}=t_{0.025,98}=1.984467"

The null hypothesis is rejected if "|t_c|\gt t_{0.025,98}."

Since "|t_c|=1.6652\lt t_{0.025,98}=1.984467," we fail to reject the null hypothesis and conclude that there is sufficient evidence to show that the difference in means between boys and girls is 3 at 5% significance level.